OpenStudy (anonymous):

Evaluate sum_{5}^{infty}(\frac{ 8 }{ 9 })^n

OpenStudy (anonymous):

@amistre64 help

hartnn (hartnn):

this related to geometric series topic, yes ? heard about it ?

OpenStudy (anonymous):

no,haven't heard about it

hartnn (hartnn):

so summation, solved any summation formula before ? \(\sum \limits_1^{\infty }a^n = \dfrac{a}{1-a} \) know this formula ?

OpenStudy (anonymous):

i got 72/9

hartnn (hartnn):

did you just use that formula ? notice that the lower limit there is +1 but the lower limit in your question is +5

OpenStudy (anonymous):

so is it a/5-a ?

OpenStudy (anonymous):

?

hartnn (hartnn):

so, you cannot use that formula directly you need to make a substitution to make the lower limit from 5 to 1 you need to put k = n-4 so that when n = 5, k =1

hartnn (hartnn):

when n= infinity, k= infinity so you get \(\sum \limits_1^{\infty }(\dfrac{8}{9})^{k+4}\) got this ? (as k= n-4, n = k+4)

hartnn (hartnn):

any doubts till here ?

OpenStudy (anonymous):

so then i get \[\sum_{1}^{5}(\frac{ 8 }{ 9 })^5\]

OpenStudy (anonymous):

?

hartnn (hartnn):

where did get that from ? no..... see we need the form a^n to use that formula , right ?

hartnn (hartnn):

so we do \((8/9)^{n+4} = (8/9)^n \times (8/9)^4\) got this ?

hartnn (hartnn):

and since the constants can be taken out of summation, we have \(\large (8/9)^4 \times \sum \limits_1^{\infty } (8/9)^n\) now you summation is of the standard form!, use the formula i gave you

hartnn (hartnn):

getting all these steps ? ask if any doubt in any step...

OpenStudy (anonymous):

ok now what?

hartnn (hartnn):

use the formula for the sum part (8/9)^4 * (....?... )

OpenStudy (anonymous):

sorry i dont get this .....i just started learning calculus .... could you explain to me step by step

hartnn (hartnn):

ok, till which step did u clearly get ?

OpenStudy (anonymous):

none

hartnn (hartnn):

did u get why i put k = n-4 ?

hartnn (hartnn):

i wanted to make the lower limit = 1 to be able to use that general formula it was 5 , so i had to subtract 4 thats why i put k = n-4 ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so thats a standard it must always be 1 ?

hartnn (hartnn):

to use that standard formula, yes so, you got till here, right ? \(\sum \limits_1^{\infty }(\dfrac{8}{9})^{k+4}\)

OpenStudy (anonymous):

yes

hartnn (hartnn):

the formula also required to have the exponent as just 'n' or 'k' but we have k+4 here so i used the exponent rule \(x^{m+n}=x^m\times x^n\)

hartnn (hartnn):

hence i got \(\large (8/9)^4 \times \sum \limits_1^{\infty } (8/9)^n\) as (8/9)^4 is constant and can be factored out

OpenStudy (anonymous):

ooo, i see it now

hartnn (hartnn):

the summation that we have now is of the standard form a^n here, a = 8/9

hartnn (hartnn):

so use that standard formula to find the value of summation ?

hartnn (hartnn):

a/(1-a) with a = 8/9

hartnn (hartnn):

once you findd the summation value, just multiply it with (8/9)^4 to get the final answer